explain domain and range of a function,in what circumstanes would the domain be something other than a real number?
Answer by richard1234(7193)

Domain = the set of inputs for a function
Range = the set of outputs for a function

The domain does not necessarily have to be a real number. For example, most functions can be defined on complex numbers; i.e. if f(x) = 5x + 2, then f(1+i) = 5(1+i) + 2 = 7 + 5i. Or, the domain of a function could be a vector quantity instead of a scalar.